OpenStudy (anonymous):

The value of an uncirculated 'Mint State - 65' 1950 Jefferson nickel minted in Denver is 8/5 the value of a 1945 nickel minted in Philadelphia in a similar condition. Together the total value of the two coins is $78. What is the value of each coin?

OpenStudy (matt101):

Let x be the value of the Philadelphia nickel. Therefore (8/5)x is the value of the Denver nickel. x + (8/5)x = 78 (13/5)x = 78 x = 30 (8/5)x = (8/5)(30) = 48 The Philadelphia nickel is worth $30, and the the Denver nickel is worth $48.

OpenStudy (anonymous):

so would philadelphia be the 1945 nickel or the 1950 nickel ?

OpenStudy (matt101):

'Mint State - 65' 1950 Jefferson nickel minted in Denver 1945 nickel minted in Philadelphia - Copied and pasted directly from your question